# Benchmarking Quantum Computers: Towards a Standard Performance Evaluation Approach

### v1

@@ II What can we learn from Classical benchmarking?
Originally, the word benchmark used to refer to ”a mark on a workbench used to compare the lengths of pieces so as to determine whether one was longer or shorter than desired” specglossary . Even though the term has clearly evolved, it retains the intuitive idea of comparing one object against another one to measure its properties. Hence, it is not a surprise that the concept of computer benchmark  [^1] We focus in this article on the study of benchmarking related to physical devices or processors, even though interesting debates are held by the community in other benchmarking fields such as algorithms benchmarking BenchmarkingOptimization ; TestingHeuristics . Indeed, similar approaches to the ones proposed for classical algorithms benchmarking could be leveraged by the quantum computing community too, for scenarios such as assessing heuristics or the overhead of implementing quantum algorithms in real devices. is based on this very same idea, and it is usually defined as ”a test, or set of tests, designed to compare the performance of one computer system against the performance of others” specglossary .
@@ II.1 Characteristics of a good benchmark
    Even when one system exhibits better performance than another on all benchmarks, the selection of an appropriate base  [^2] In a ratio game, we are comparing devices performance using ratios, where the denominator is called the base. can be manipulated to showcase a more favorable relative performance.
@@ III Quantum benchmarking
Quantum computers are devices that leverage the principles of quantum mechanics to address challenges beyond the computational capacity of classical computers. Several physical implementations for quantum computers have been explored, each employing different technologies as qubits, such as trapped ion QCion ; Zhang_2017 , superconducting circuits Arute2019 ; ZHU2022240 ; PhysRevLett.127.180501 , photonic QCPhoton ; PhysRevLett.127.180502 , and cold atoms Bernien2017ProbingMD ; Ebadi2020QuantumPO . Each of these architectures has its own set of advantages and disadvantages, and research is ongoing to determine which approach is most suitable for scaling up quantum computers to a practical level. Additionally, various quantum computing paradigms have emerged, including quantum annealing for optimization problems; analog quantum computing, in which a controllable quantum device used to simulate other quantum systems; and digital quantum computing, which uses a gate-based quantum devices to simulate a quantum system. These three paradigms offer different perspectives on the practical applications of quantum computing.
@@
The diversity of technological platforms and paradigms used in the development of quantum computers introduces significant complexity in comparing the performance among different devices. For this reason, the search for a common benchmark that allows comparison of quantum computers performance across different technological platforms and paradigms poses a challenge to the scientific community. Despite these challenges, efforts have been made to develop benchmarks enabling comparison among diverse quantum computers. Primarily, these efforts focus on three types of relevant comparisons: comparing the performance of different quantum processors within the same computational paradigm (e.g., comparison of digital quantum computing platform providers such as Forest, Qiskit, ProjectQ LaRose2019overviewcomparison , or comparisons of trapped atomic ions and superconducting circuits Linke2017 ); comparing different computational paradigms (e.g., between processors for digital quantum computing and quantum annealing or digital-analog quantum computing Flannigan_2022 ); and finally, comparing quantum computers to classical computers (e.g. comparing digital processing against quantum annealing Jnger2021 , empirical comparisons relying upon simplified Machine Learning (ML) tasks havenstein2018comparisons ; Cugini2023 or even the claims on quantum supremacy Arute2019 ; kalai2022google ). It is worth noting that these three comparisons are crucial for advancing scalable and reliable quantum technologies, but each of them demands a thorough analysis. Furthermore, a similar analysis could be expanded to include algorithmic comparisons. Looking ahead, benchmarking should be extended to cover other potential new aspects of a quantum computer that we have not even envisioned yet (e.g. qRAM QuantumRam or other future components that we have not even thought about yet). During this work, we will restrict ourselves to the evaluation of the performance of quantum processors.
@@ III.1 Nuances of benchmarking quantum computers
  • A quantum computer execution typically sticks to the Input-Process-Output (IPO) model, a rather general framework used in software engineering to describe the structure of an information processing program. Generally, a quantum computer needs to complete three different stages: the input or loading of information, the processing thereof, and the output or retrieval of relevant information. Most quantum algorithms adhere to this model  [^3] There are notable exceptions, such as hybrid classical-quantum algorithms., where the input is state preparation or matrix exponentiation, and the output is just the quantum measurement (see Fig.1), adding extra steps to the classical algoritmic execution.
@@ III.2 A common language to discuss
The underlying goal of a quantum benchmark is to be able to compare devices with each other. What is more, similarly to classical benchmarking, quantum benchmarks involve running specific tests or workloads on the systems and software collecting data to analyze its performance. QASMBench li2022qasmbench , Quantum Volume PhysRevA.100.032328 , or Quantum LINPACK QuantumLINPACK are examples of quantum benchmarks.
When executing these benchmarks, we need to later compile a list of metrics that will describe the system, we introduce the definition of a performance metric analogous to the classical case:
@@ III.3 Reviewing existing proposals for benchmarking
Recently, there have been attempts to categorize benchmarking practices. However, the lack of a precise definition of concepts has led to confusion in the field of quantum benchmarking. There are instances where the concepts of certification and benchmarking are intertwined within a shared framework eisert2020quantum , or concepts such as metrics and benchmarks are sometimes used interchangeably SoKBenchmarkingQuantumComputer . In Ref. SoKBenchmarkingQuantumComputer , the term physical benchmarkings has been introduced to describe properties of a quantum processor that are more accurately understood as metrics. Additionally, the concept of aggregated benchmarks has been presented, which combines metrics and benchmarking techniques. For instance, Circuit Layer Operations per Second https://doi.org/10.48550/arxiv.2110.14108 is a metric while Mirroring Circuits Proctor2021 ) is a technique employed in benchmarks. Other proposals more aligned with the intent of distinguishing between different terms have been presented, although their discussion has been limited. An excellent example of such studies is the article by Resch and Karpuzcu on the impact of quantum noise https://doi.org/10.48550/arxiv.1912.00546 .

### v2

@@ II What can we learn from Classical benchmarking?
Originally, the word benchmark used to refer to ”a mark on a workbench used to compare the lengths of pieces so as to determine whether one was longer or shorter than desired” specglossary . Even though the term has clearly evolved, it retains the intuitive idea of comparing one object against another one to measure its properties. Hence, it is not a surprise that the concept of computer benchmark PP is based on this very same idea, and it is usually defined as ”a test, or set of tests, designed to compare the performance of one computer system against the performance of others” specglossary .
@@ II.1 Characteristics of a good benchmark
    Even when one system exhibits better performance than another on all benchmarks, the selection of an appropriate base P can be manipulated to showcase a more favorable relative performance.
@@ III Quantum benchmarking
Quantum computers are devices that leverage the principles of quantum mechanics to address challenges beyond the computational capacity of classical computers. Several physical implementations for quantum computers have been explored, each employing different technologies as qubits, such as trapped ion QCion ; Zhang_2017 , superconducting circuits Arute2019 ; ZHU2022240 ; PhysRevLett.127.180501 , photonic QCPhoton ; PhysRevLett.127.180502 , and cold atoms Bernien2017ProbingMD ; Ebadi2020QuantumPO . Each of these architectures has its own set of advantages and disadvantages, and research is ongoing to determine which approach is most suitable for scaling up quantum computers to a practical level. Additionally, various quantum computing paradigms have emerged, including quantum annealing for optimization problems Rajak_2022 ; analog quantum computing, in which a controllable quantum device used to simulate other quantum systems Kendon_2010 ; and digital quantum computing, which uses a gate-based quantum devices to simulate a quantum system doi:10.1126/science.273.5278.1073 . These three paradigms offer different perspectives on the practical applications of quantum computing.
@@
The diversity of technological platforms and paradigms used in the development of quantum computers introduces significant complexity in comparing the performance among different devices. For this reason, the search for a common benchmark that allows comparison of quantum computers performance across different technological platforms and paradigms poses a challenge to the scientific community. Despite these challenges, efforts have been made to develop benchmarks enabling comparison among diverse quantum computers. Primarily, these efforts focus on three types of relevant comparisons: comparing the performance of different quantum processors within the same computational paradigm (e.g., comparison of digital quantum computing platform providers such as Forest, Qiskit, ProjectQ LaRose2019overviewcomparison , or comparisons between quantum processors Linke2017 ; miessen2024benchmarkingdigitalquantumsimulations ; Ilin_2024 ; wright2024noisyintermediatescalequantumsimulation ); comparing different computational paradigms (e.g., between processors for digital quantum computing and quantum annealing or digital-analog quantum computing Flannigan_2022 ); and finally, comparing quantum computers to classical computers (e.g. comparing digital processing against quantum annealing Jnger2021 , empirical comparisons relying upon simplified Machine Learning (ML) tasks havenstein2018comparisons ; Cugini2023 or even the claims on quantum supremacy Arute2019 ; kalai2022google ). It is worth noting that these three comparisons are crucial for advancing scalable and reliable quantum technologies, but each of them demands a thorough analysis. Furthermore, a similar analysis could be expanded to include algorithmic comparisons. Looking ahead, benchmarking should be extended to cover other potential new aspects of a quantum computer that we have not even envisioned yet (e.g. qRAM QuantumRam or other future components that we have not even thought about yet). During this work, we will restrict ourselves to the evaluation of the performance of quantum processors.
@@ III.1 Nuances of benchmarking quantum computers
  • A quantum computer execution typically sticks to the Input-Process-Output (IPO) model, a rather general framework used in software engineering to describe the structure of an information processing program. Generally, a quantum computer needs to complete three different stages: the input or loading of information, the processing thereof, and the output or retrieval of relevant information. Most quantum algorithms adhere to this model PPP , where the input is state preparation or matrix exponentiation, and the output is just the quantum measurement (see Fig.1), adding extra steps to the classical algoritmic execution.
@@ III.2 A common language to discuss
The underlying goal of a quantum benchmark is to be able to compare devices with each other. What is more, similarly to classical benchmarking, quantum benchmarks involve running specific tests or workloads on the systems and software collecting data to analyze its performance. Examples of quantum benchmarks include QASMBench li2022qasmbench , Quantum Volume PhysRevA.100.032328 , or Quantum LINPACK QuantumLINPACK .
When executing these benchmarks, we need to later compile a list of metrics that will describe the system. We introduce the definition of a performance metric analogous to the classical case:
@@ III.3 Reviewing existing proposals for benchmarking
Recently, there have been attempts to categorize benchmarking practices. However, the lack of a precise definition of concepts has led to confusion in the field of quantum benchmarking. There are instances where the concepts of certification and benchmarking are intertwined within a shared framework eisert2020quantum , or concepts such as metrics and benchmarks are sometimes used interchangeably SoKBenchmarkingQuantumComputer . In Ref. SoKBenchmarkingQuantumComputer , the term physical benchmarkings has been introduced to describe properties of a quantum processor that are more accurately understood as metrics. Additionally, the concept of aggregated benchmarks has been presented, which combines metrics and benchmarking techniques. For instance, Circuit Layer Operations per Second https://doi.org/10.48550/arxiv.2110.14108 is a metric while Mirroring Circuits Proctor2021 ) is a technique employed in benchmarks. Other proposal more aligned with the intent of distinguishing between different terms have been presented proctor2024benchmarkingquantumcomputers . An excellent example of such studies is the article by Resch and Karpuzcu on the impact of quantum noise https://doi.org/10.48550/arxiv.1912.00546 .